OpenStudy (anonymous):

solve the systems of equations by graphing: 3x-3y=0 5x-5y=0

OpenStudy (nikita2):

If you try to draw those graph you'll see that they are same.

OpenStudy (nikita2):

For example, if point A (a1,a2) belongs to first one, then A belongs to second too. Because 3*a1 - 3*a2 =0 => a1 - a2 =0 => 5*a1 - 5*a2 =0